OpenStudy (anonymous):

I need help on this problem. I looked at an example but it it really confusing. Solve the triangle.... a=1,b=4,C=60° c= A= B=

OpenStudy (anonymous):



OpenStudy (anonymous):

cos theorem` c^2=a^2+b^2-2*a*b*cos(C) so c^2=1+16-8*1/2 c^2=13 c=sqrt(13)

OpenStudy (anonymous):

then, sin thorem a/sin(A)=b/sin(B)=c/sin(C)

OpenStudy (anonymous):

so that is to find C, and square root of 13 is 169

OpenStudy (anonymous):

nor square of 13 is 169, square root of 13 is eqaul to about 3.6, but you can write it sqrt(13)

OpenStudy (anonymous):

thats right, 3.6, so next part is what I get stuck at, cos^-1(1+13-4/4*1*square root of 13)

OpenStudy (anonymous):

no you just need to solve it according to sin theorem it is a/sin(A)=b/sin(B)=c/sin(C) 1/sin(A)=sqrt(13)/sin(60) then sin(A)=sin(60)/sqrt(13) and sin(B)=4*sin(60)/sqrt(13)

OpenStudy (anonymous):

ok, is that to find A?

OpenStudy (anonymous):

I am confused how to figure that part out?

OpenStudy (anonymous):

if you found sin(A), that means you also found A, because then arcsin(sin(A))

OpenStudy (anonymous):

ok, so to find A do I plug in what you wrote earlier?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin(60)=sqrt(3)/2 so sin(A)=sqrt(3)/(2*sqrt(13) A=arcsin(sqrt(3)/(2*sqrt(13))